Status Report

SATOPS Morning Report 12 Feb 2001

By SpaceRef Editor
February 12, 2001
Filed under ,

SATOPS MORNING REPORT February 12, 2001

 

POES

No change in the status of the POES spacecraft.

NOAA-16 rev 1992 / F at 1617z on Friday, February 9 – Completed dumping
the final 16K of memory from CPU2 (standard ground image maintenance).

NOAA-14 rev 31519 / F at 0301z on Saturday, February 10 – GAC5B received
noisy from spacecraft with 93% of the expected data received. Recorder
management prevented replay from spacecraft.

NOAA-14 rev 31525 / F at 1329z on Saturday, February 10 – LAC5B received
noisy from spacecraft with 96.7% of the expected data received. No re-ingest
required.

NOAA-15 rev 14296 / F at 1953z on Sunday, February 11 – GAC5B received
short from spacecraft with 95.8% of the expected data received.  No
re-ingest required.

NOAA-12 rev 50634 / F at 2125z on Sunday, February 11 – GAC3A was lost
due to no frame synch lock (not recoverable from IMUX system). Recorder
management prevented replay from spacecraft.

POES operations were nominal over the past 72 hours.

 

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 72 hours.

GOES

No change in the status of the GOES spacecraft.

Battery reconditioning efforts continue on all spacecraft.

At 1733z on February 9, a GOES-8 frame break was observed and attributed
to loss of signal synch.

At 1435z on February 9, GOES-10 frame breaks were observed and attributed
to command I/O error.

At 1646z on February 9, GOES-10 frame breaks were observed and attributed
to loss of signal synch.

At 1550z on February 9, the GOES-10 Imager patch temperature was changed
from the Med to Low set point.  CDA was directed to set SPS3 &
SPS4 gain 1 cal mode set to level 1, as per the procedure.

GOES operations were nominal over the past 72 hours.

 

 

Questions or comments to msimpson@nesdis.noaa.gov

SpaceRef staff editor.